OpenStudy (anonymous):

f(x)=|x| g(x)=|x-3| solve : f+g ?? guys help

OpenStudy (tkhunny):

(f+g)(x) = |x|+|x-3| ?? Would it be something else? Perhaps you are trying to simplify it, or something?

OpenStudy (anonymous):

yeah I need that

OpenStudy (tkhunny):

??

OpenStudy (anonymous):

I need to solve this |x|+|x-3|

OpenStudy (anonymous):

how

OpenStudy (tkhunny):

There is nothing to solve, there. Do you mean |x|+|x-3| = 0, or perhaps equal to something else? Perhaps we're just evaluating it at x = 0 or x = 3? "Solve" just doesn't make any sense without an equation of some sort. Do we have the entire problem statement?

OpenStudy (anonymous):

Evaluate the function at the indicated values f(x)=|x| g(x)=|x-3| f+g and f-g and f*g and f/g and g/f

OpenStudy (anonymous):

@tkhunny that's all the question is

OpenStudy (tkhunny):

Aha!! "At the indicated values". That's what we were missing. There must be some values indicated. x = 0? (f+g)(x) = |x| + |x-3| (f-g)(x) = |x| - |x-3| (f*g)(x) = |x| * |x-3| (f/g)(x) = |x| / |x-3| (g/f)(x) = |x-3| / |x| Now, we need some "indicated points".

OpenStudy (anonymous):

ok can you explain more I didn't get it sorry

OpenStudy (anonymous):

@hartnn sir help

OpenStudy (tkhunny):

We still need "indicated points". Let's use x = 0 for an example. (f+g)(0) = |0| + |0-3| = 0 + |-3| = 0+3 = 3 (f-g)(0) = |0| - |0-3| = 0 - |-3| = 0-3 = -3 (f*g)(0) = |0| * |0-3| = 0*|-3| = 0 (f/g)(0) = |0| / |0-3| = 0/|-3| = 0 (g/f)(0) = |0-3| / |0| = |-3|/0 -- Whoops!! Have to let this one go. Can't have a zero down there in the denominator.